### v1

# Prominent bump in the two-neutron separation energies of neutron-rich lanthanum isotopes revealed by high-precision mass spectrometry

Abstract
We report on high-precision atomic mass measurements of ${}^{148\text{-}153}$a and 151Ce performed with the JYFLTRAP double Penning trap using the Phase-Imaging Ion-Cyclotron-Resonance technique. The masses of 152,153La were experimentally determined for the first time. We confirm the sharp kink in the two-neutron separation energies at the neutron number ${N=93}$ in the cerium (${Z=58}$) isotopic chain. Our precision mass measurements of the most exotic neutron-rich lanthanum (${Z=57}$) isotopes reveal a sudden increase in two-neutron separation energies from ${N=92}$ to ${N=93}$. Unlike in the cerium isotopic chain, the kink is not sharp but extends to ${N=94}$ forming a prominent bump. The gain in energy is about 0.4 MeV, making it one of the strongest changes in two-neutron separation energies over the whole chart of nuclides, away from nuclear shell closures. The results call for further studies to elucidate the structure of neutron-rich lanthanum isotopes.
Penning-trap mass spectrometry can probe one of the fundamental properties of a nucleus, its mass, with high precision [1, 2]. Nuclear binding energies $BE$ are directly related to atomic masses, and provide information on, e.g., interactions between nucleons (neutrons and protons), nuclear shell structure and deformation (see e.g. Refs. [3, 4]). Even subtle changes in nuclear structure can be probed via differences in binding energies along an isotopic chain, when observing abrupt deviations from the otherwise rather linear trend of the two-neutron separation energies $S_{2n}=BE(Z,N)-BE(Z,N-2)$, where $Z$ and $N$ denote the number of protons and neutrons, respectively.
The most prominent example observed throughout the entire chart of nuclides is a sharp drop of the $S_{2n}$ values after crossing a magic neutron number, associated with a nuclear shell closure. Outside closed neutron shells, notable changes in the $S_{2n}$ trends have only been identified in three regions, located at ${N\approx 32}$, ${N\approx 60}$ and ${N\approx 90}$. In the first region, the high-precision mass measurements of K [5], Ca [6, 7], Sc [8, 9], Ti and V [9] isotopes at TITAN, ISOLTRAP and BigRIPS-SLOWRI showed a sudden decrease in the $S_{2n}$ values, by about 3 MeV, between ${N=32}$ and ${N=34}$. This behavior indicates a strong shell closure at ${N=32}$ and it might be interpreted as an emergence of a new magic number [4]. At the same time, the results of laser-spectroscopy studies contradict this interpretation [10, 11].
In the $N\approx 60$ region, onset of deformation was predicted [12, 13] and experimentally discovered [14] already in 1970. Since then, it has been intensively studied by $\gamma$-, conversion-electron- and laser-spectroscopy, as well as mass measurements, see e.g. Ref. [15] and references therein. The mass measurements in the neutron-rich isotopic chains between ${Z=37}$ and ${Z=42}$ [16, 17, 18, 19, 20, 21] revealed a strong kink in the $S_{2n}$ values at $N\approx 60$, deviating from a linear trend by up to 2 MeV. This change has been associated with shape coexistence [15], deformed and spherical orbitals crossing each other [22], and quantum phase transition [23].
In the neutron-rich rare-earth region, the onset of strong quadrupole deformation at $N\approx 90$ was discovered already in the 1950s and associated with the breaking up of the $h_{11/2}$ shell [24, 25, 26]. This change in nuclear structure is observed as a kink or a broader bump in the $S_{2n}$ values for isotopic chains from praseodymium (${Z=59}$) to dysprosium (${Z=66}$), increasing the $S_{2n}$ values by up to 1.5 MeV [27]. The kinks have been later interpreted as being due to first-stage quantum shape phase transition from spherical to prolate deformation [28]. However, recent calculations have indicated that the structure may not be as simple as traditionally predicted but can also include triaxiality, shape coexistence and octupolar deformation (see e.g. Ref. [29]).
The masses of neutron-rich isotopes ranging from the barium ($Z=56$) to holmium ($Z=67$) isotopic chains have been extensively studied with the JYFLTRAP Penning trap [30, 31] and the Canadian Penning trap (CPT) [32, 33, 34, 35]. In the case of the neutron-rich barium isotopes, a relatively flat trend in the $S_{2n}$ values is observed. While the lanthanum (${Z=57}$) and cerium (${Z=58}$) isotopic chains exhibit a small kink in the $S_{2n}$ values at $N=93$, the lack of precise experimental data hinders drawing clear conclusions on possible structural effects. The masses of 149,151La are known with a limited precision of 200 and 440 keV [36, 37, 27], respectively, and the mass of 150La is known from a high-precision mass measurement performed at CPT [34]. In the case of 151Ce93 isotope, the deviation from a smooth trend could be explained by a presence of a long-lived isomer proposed in the ENSDF evaluation [38] but not observed in Ref. [32].
In addition to poorly known mass values, the understanding of possible structural changes in the lanthanum isotopic chain around ${N\approx 93}$ is hindered by the scarcity of spectroscopic information [39]. The most neutron-rich lanthanum isotope studied via $\gamma$-ray spectroscopy to date remains 149La (${N=92}$) where a strong quadrupole deformation is predicted to develop [40, 41]. The $\gamma$-ray spectroscopy on 145La88 and 147La90 have shown evidence of octupole correlations [42, 43, 44, 45], similarly to the neighboring even-$Z$ chains of barium [46, 47, 48, 49, 50, 51, 52, 53, 54] and cerium [53, 54, 55, 56, 57], where strong indications of octupole deformation and correlations have been observed and predicted for nuclei around 146Ba by density functional theory (DFT) calculations [58].
In this Letter, we report on the precise mass measurements of neutron-rich 148-153La and 151Ce isotopes performed with the JYFLTRAP double Penning-trap mass spectrometer [59, 60] using the Phase-Imaging Ion-Cyclotron-Resonance (PI-ICR) method [61, 62, 63, 64]. Our work unveils a prominent bump in the $S_{2n}$ values at ${N\approx 92-94}$ for the lanthanum isotopic chain, indicating a sudden change in the nuclear structure. We also confirm the sharp kink in the $S_{2n}$ values at $N=93$ for the cerium isotopes.
The measurement was performed at the Ion Guide Isotope Separator On-Line (IGISOL) facility [65, 66]. The neutron-rich rare-earth isotopes were produced using $25$-MeV protons impinging on a $15$-mg/cm2-thick ${}^{\text{nat}}$ target. The fission fragments, mostly singly charged, were stopped in a helium-filled gas cell (at about 300 mbar) before extracted out through a sextupole ion guide [67]. The ions were further accelerated to $30q$ kV and separated based on their mass-to-charge ratio $m/q$ using a $55$ degree dipole magnet. The continuous beam was cooled and bunched with the Radiofrequency Quadrupole Cooler-Buncher [68] and sent to JYFLTRAP.
Figure 1: Phase projection of a) the magnetron and b) cyclotron motion of 152La+ ions on the position-sensitive detector with the PI-ICR technique using $t_{\text{acc}}=92$ ms. The polar angles of the magnetron ($\alpha_{-}$) and cyclotron ($\alpha_{+}$) phase spots are indicated with an arc. The average cyclotron excitation radius is marked with the dashed circle on b).
The mass-selective buffer-gas cooling technique [69] was applied in the first trap of JYFLTRAP to select the ions of interest for the mass measurements, which were performed with the PI-ICR method [61, 62, 63, 64] in the second trap. The ion’s free space cyclotron frequency $\nu_{c}=qB/(2\pi m)$, where $B$ is the magnetic field strength in the trap, was determined based on the detected radial eigenmotion phase spots on a position-sensitive detector after a chosen accumulation time $t_{\text{acc}}$ and $n_{c}$ revolutions [61, 62, 63, 64], following $\nu_{c}=[(\alpha_{+}-\alpha_{-})+2\pi n_{c}]/(2\pi t_{acc})$, where $\alpha_{\pm}$ are the polar angles of the cyclotron ($+$) and magnetron (--) phase images with respect to the trap center on the detector. A typical PI-ICR figure obtained in this work is shown on Fig. 1.
The magnetic field strength $B$ was determined using 133Cs as a reference. The atomic mass was determined from the frequency ratio between the reference ions and ions of interest (both singly charged), $r=\nu_{c\text{,ref}}/\nu_{c}$, as $m=r(m_{\text{ref}}-m_{e})+m_{e}$, where $m_{e}$ is the free-electron mass. The number of detected ions per bunch was limited to up to one to reduce the effects coming from ion-ion interactions [70]. Systematic effects due to the temporal fluctuations of the magnetic field, mass dependency, and the distortion of the ions’ motion projection, were taken into account, together with a residual uncertainty. The contribution of these effects is detailed in Ref. [64]. A thorough description of the experimental methods used in this work can be found in Ref. [31].
[…]
Nuclide $T_{1/2}$ $J^{\pi}$ $t_{acc}$ (ms) $r=\nu_{\text{c,ref}}/\nu_{\text{c,ion}}$ $\text{ME}_{\text{JYFLTRAP}}$ (keV) $\text{ME}_{\text{AME20}}$ (keV) Diff. (keV)
${}^{148}\text{La}$ $1.414(25)$ s (2-) 551 1.113 067 600(22)uncertain1.113067600221.113\,067\,600(22)start_ARG 1.113 067 600 end_ARG start_ARG ( 22 ) end_ARG 62 698.5(28)uncertain-62698.528-62\,698.5(28)start_ARG - 62 698.5 end_ARG start_ARG ( 28 ) end_ARG 62 709(19)uncertain-6270919-62\,709(19)start_ARG - 62 709 end_ARG start_ARG ( 19 ) end_ARG 11(19)
${}^{149}\text{La}$ $1.071(22)$ s (3/2-) 344 1.120 613 668(23)uncertain1.120613668231.120\,613\,668(23)start_ARG 1.120 613 668 end_ARG start_ARG ( 23 ) end_ARG 59 988.3(28)uncertain-59988.328-59\,988.3(28)start_ARG - 59 988.3 end_ARG start_ARG ( 28 ) end_ARG 60 220(200)uncertain-60220200-60\,220(200)start_ARG - 60 220 end_ARG start_ARG ( 200 ) end_ARG 232(200)
${}^{150}\text{La}$ $504(15)$ ms (3+) 313 1.128 167 562(28)uncertain1.128167562281.128\,167\,562(28)start_ARG 1.128 167 562 end_ARG start_ARG ( 28 ) end_ARG 56 309.4(35)uncertain-56309.435-56\,309.4(35)start_ARG - 56 309.4 end_ARG start_ARG ( 35 ) end_ARG 56 311.1(25)uncertain-56311.125-56\,311.1(25)start_ARG - 56 311.1 end_ARG start_ARG ( 25 ) end_ARG 2(4)
${}^{151}\text{La}$ $465(24)$ ms ${}^{+}\#$ 210 1.135 714 229(23)uncertain1.135714229231.135\,714\,229(23)start_ARG 1.135 714 229 end_ARG start_ARG ( 23 ) end_ARG 53 524.5(29)uncertain-53524.529-53\,524.5(29)start_ARG - 53 524.5 end_ARG start_ARG ( 29 ) end_ARG 53 310(440)uncertain-53310440-53\,310(440)start_ARG - 53 310 end_ARG start_ARG ( 440 ) end_ARG [^1] Also 53 542(17)uncertain-5354217-53\,542(17)start_ARG - 53 542 end_ARG start_ARG ( 17 ) end_ARG keV in Ref. [72]. $-215(440)$
${}^{152}\text{La}$ $287(16)$ ms ${}^{-}\#$ 92 1.143 273 524(85)uncertain1.143273524851.143\,273\,524(85)start_ARG 1.143 273 524 end_ARG start_ARG ( 85 ) end_ARG 49 178(11)uncertain-4917811-49\,178(11)start_ARG - 49 178 end_ARG start_ARG ( 11 ) end_ARG 49 290(300)uncertain-49290300-49\,290(300)start_ARG - 49 290 end_ARG start_ARG ( 300 ) end_ARG##\## 112(300)##\##
${}^{153}\text{La}$ $245(18)$ ms ${}^{+}\#$ 83 1.150 823 06(30)uncertain1.15082306301.150\,823\,06(30)start_ARG 1.150 823 06 end_ARG start_ARG ( 30 ) end_ARG 46 038(37)uncertain-4603837-46\,038(37)start_ARG - 46 038 end_ARG start_ARG ( 37 ) end_ARG
177 1.150 823 12(15)uncertain1.15082312151.150\,823\,12(15)start_ARG 1.150 823 12 end_ARG start_ARG ( 15 ) end_ARG 46 031(18)uncertain-4603118-46\,031(18)start_ARG - 46 031 end_ARG start_ARG ( 18 ) end_ARG
Weighted average: 46 032(16)uncertain-4603216-46\,032(16)start_ARG - 46 032 end_ARG start_ARG ( 16 ) end_ARG 46 060(300)uncertain-46060300-46\,060(300)start_ARG - 46 060 end_ARG start_ARG ( 300 ) end_ARG##\## 28(300)##\##
${}^{151}\text{Ce}$ $1.76(6)$ s (3/2-) 335 1.135 652 075(22)uncertain1.135652075221.135\,652\,075(22)start_ARG 1.135 652 075 end_ARG start_ARG ( 22 ) end_ARG 61 219.7(27)uncertain-61219.727-61\,219.7(27)start_ARG - 61 219.7 end_ARG start_ARG ( 27 ) end_ARG 61 225(18)uncertain-6122518-61\,225(18)start_ARG - 61 225 end_ARG start_ARG ( 18 ) end_ARG [^2] Also 61 230(15)uncertain-6123015-61\,230(15)start_ARG - 61 230 end_ARG start_ARG ( 15 ) end_ARG keV in Ref. [72]. 5(18)
The determined frequency ratios $r$ and the corresponding mass-excess (ME) values are reported in Table 1, together with the comparison to the Atomic Mass Evaluation 2020 (AME20) [27]. The masses of 152,153La were experimentally determined for the first time and are in agreement with the AME20 extrapolations based on systematics. With the PI-ICR technique, the precision for all studied nuclides was greatly improved. The results are in agreement with the previously known mass values [37, 32, 34] except for 149La, which deviates by more than $1\sigma$ from the value based on a $\beta$-decay study [36]. We observed only one state in 151Ce, similarly to the CPT measurement [32], thus not supporting the existence of a long-lived isomer proposed in the ENSDF evaluation [38].
Figure 2: Two-neutron separation energies $S_{2n}$ based on the new JYFLTRAP mass measurements from this work (in red) in comparison with the experimental AME20 values [27] (in black). For the cerium (Ce), praseodymium (Pr) and promethium (Pm) isotopic chains, also the recent CPT values are shown [35] (in blue).
Figure 3: Two-neutron separation energies $S_{2n}$ determined from this work (in red), together with the $S_{2n}$ calculated from the AME20 experimental mass values [27] (in black) and various mass models, namely DZ [73], WS4 [74], FRDM2012 [75], UNEDF1 [76], SKP [77] and BSkG3 [78].
Figure 4: Single-particle diagrams adapted from Ref. [79] for a) protons around $Z=58$, b) neutrons around $N=88$, and c) neutrons around $N=60$. The levels are plotted with the deformation parameter $\delta(\approx 0.95\beta)$ as defined in [80]. Even and odd parity are shown with solid and dashed lines, respectively. In all three regions, the upward-sloping extruder orbital crosses the downward-sloping deformation-driving orbitals. For the $N=88$ and $N=60$ regions, this has served an explanation for the observed shape coexistence and onset of ground-state deformation (see e.g. Ref. [15, 23, 29, 81].
Figure 2 shows two-neutron separation energies for the isotopic chains from caesium (${Z=55}$) to promethium (${Z=61}$) based on AME20 [27], including our new results on 148-153La and 151Ce, as well as the recent CPT measurements for cerium, praseodymium and promethium isotopic chains [35]. The values were obtained from the mass-excesses as $S_{2n}(Z,N)=\text{ME}(Z,N-2)-\text{ME}(Z,N)+2\text{ME}_{n}$, where $\text{ME}_{n}$ is the neutron mass excess. While the heavier isotopic chains in the rare-earth region exhibit prominent kinks at $N\approx 90$ due to onset of strong quadrupole deformation, the effect seems to vanish when moving to lower proton numbers according to AME20 values. This is in accordance with the energies of the first $2^{+}$ states, which decrease steeply at $N\geq 88$ for the even-even nuclei in the region but much less so for the lower-$Z$ isotopic chains of cerium and barium, which reach the minimum at larger $N$ [39].
In this work, we provide new, precise $S_{2n}$ values for neutron-rich lanthanum isotopes and extend the known values up to 153La (see Fig. 2). These reveal a pronounced $\approx$0.4 MeV bump in the $S_{2n}$ values at neutron numbers $N=93-94$. Moreover, we can confirm the sharp small kink in the cerium $S_{2n}$ values at $N=93$ with the measurement of 151Ce.
The experimental $S_{2n}$ values of lanthanum isotopes are compared to various mass models in Fig. 3. None of the models, which include both phenomenological macroscopic-microscopic and self-consistent mean-field approaches [82] based on energy density functionals (EDFs), can reproduce the bump observed at $N=93-94$ but remain compatible with a rather smooth linear decrease in the $S_{2n}$ values. It should be noted that the mass models have typical root-mean-square uncertainties of a few hundred keVs and have challenges to reproduce the observed kinks in the ${N\approx 60}$ and ${N\approx 90}$ regions (see e.g. Refs. [30, 83]). Moreover, the odd-$A$ and odd-odd nuclei are not explicitly calculated in EDF-based mass models but estimated from the pairing gap of neighboring even-even nucleus, or if calculated at the Hartree-Fock-Bogoliubov level using the quasiparticle blocking procedure, various polarization effects may appear and affect the predictions [82]. Thus, the situation presented in Fig. 3 is not unusual.
The $N\approx 60$ and $N\approx 90$ regions have been extensively studied and the prominent kinks in the $S_{2n}$ values observed in multiple isotopic chains (see e.g. Refs. [15, 84] and references therein). At the same time, in the case of the lanthanum isotopes, our mass measurements provide the first experimental data indicating structural changes beyond ${N=92}$ and suggest that the bump at ${N=93-94}$ in the $S_{2n}$ values is a unique feature of this one element, accompanied with a smaller sharp kink for the cerium isotopes at $N=93$. While definite conclusions on the nature of the structural change cannot be drawn based on masses, spectroscopic studies can shed light on it.
The region of neutron-rich barium to cerium isotopes is complex in structure, including both octupole and quadrupole deformations and shape coexistence. An island of octupole deformation centered around 146Ba [85, 86] expands to neighbouring lanthanum and cerium chains. For the latter, the spectroscopic studies suggest a gradual transition from octupolar correlations in 150,152Ce [87, 88] to a rigid rotor at 154Ce [89]. The recent study of neutron-rich barium isotopes from the ISOLDE Decay Station indicates an increasing collectivity in 148-150Ba towards prolate deformation [90]. The findings of these shape changes are further supported by theoretical calculations for the barium [90] and cerium [81] chains, both indicating octupolar deformation for lighter isotopes and shape coexistence, evolving toward prolate ground-state deformation at heavier isotopes.
For the lanthanum isotopes, the spectroscopic studies of the most exotic isotope studied so far, 149La [40, 41], propose a transition from octupolar to prolate deformation at ${N\geq 92}$. This coincides with the onset of the bump in the two-neutron separation energies, see Fig. 2. Urban et al. [89] have discussed the potential role of the ${9/2^{+}[404]}$ proton extruder orbital in creating deformation in the neutron-rich ${A\approx 150}$ region. The ${9/2^{+}[404]}$ proton extruder brings in two additional protons to the Fermi level that can be further transferred to the $3/2^{-}[541]$ deformation-driving down-sloping orbital (see Fig. 4a) [91]. The mechanism would be analogous to the neutron extruder orbitals ${11/2^{-}[505]}$ and ${9/2^{+}[404]}$ (see Figs. 4b and c), which bring neutrons to the Fermi surface that can be transferred to deformation-driving intruder orbitals.
While the neutron extruder orbitals and their role in the onset of deformation observed at ${N=88-90}$ [25] and ${N\approx 60}$ [91] has been highlighted in the past, the role of the ${9/2^{+}[404]}$ proton extruder orbital has remained inconclusive. The prominent bump in the two-neutron separation energies of the lanthanum isotopes at $N=93-94$ and the smaller, sharp kink for the cerium isotopes at $N=93$ serve potential evidence for such a phenomenon. The location of the bump at a higher neutron number than in the heavier isotopic chains might also be explained by the proton extruder orbital. Namely, proton-neutron interactions play a crucial role in creating deformation which correlates with the number of valence protons and neutrons [92, 93]. The proton extruder brings in two more valence protons, requiring two more neutrons to reach the same number of valence protons and neutrons as compared to the heavier isotopic chains [89].
In this Letter we have reported on the high-precision mass measurements of 148-153La and 151Ce performed using the PI-ICR technique with the JYFLTRAP double Penning trap. With the new mass data, we could pin down the trends in the two-neutron separation energies $S_{2n}$. Our value for 151Ce confirms the small peak at ${N=93}$ in the cerium chain, coinciding with the transitional region from octupolar to quadrupolar deformation previously studied via $\gamma$-ray spectroscopy. For the lanthanum isotopes, our new values reveal a bump about 0.4-MeV in the $S_{2n}$ values at neutron numbers ${N=92-94}$, a unique feature within the neighboring isotopic chains which none of the studied mass models could reproduce. The change in the $S_{2n}$ trend agrees with the predicted onset of quadrupole deformation at ${N\geq 92}$ that could be potentially caused by the ${9/2^{+}[404]}$ proton extruder orbital. Such a proton extruder could be a plausible explanation for the localisation of the bump within the lanthanum and cerium isotopic chains, as for the other well-known kinks in two-neutron separation energies that span over many isotopic chains, neutron extruder orbitals play the main role. More extensive studies on exotic neutron-rich lanthanum isotopes at ${N\geq 90}$ would be highly welcome to elucidate the evolution of nuclear structure in the region.
Acknowledgements.
We acknowledge the funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 771036 (ERC CoG MAIDEN), No. 101057511 (EURO-LABS) and No. 861198–LISA–H2020-MSCA-ITN-2019 and from the Research Council of Finland projects No. 354589, No. 295207, No. 306980, No. 327629 and No. 354968. The financial support provided by the Vilho, Yrjö and Kalle Väisälä Foundation is acknowledged by J. Ruotsalainen. We thank Markus Kortelainen for insightful discussions.

### v2

@@
 We confirm the sharp kink in the two-neutron separation energies at the neutron number ${N=93}$ in the cerium (${Z=58}$) isotopic chain.
-Our precision mass measurements of the most exotic neutron-rich lanthanum (${Z=57}$) isotopes reveal a sudden increase in two-neutron separation energies from ${N=92}$ to ${N=93}$.
+Our precision mass measurements of the most exotic neutron-rich lanthanum (${Z=57}$) isotopes reveal a unexpected sudden increase in two-neutron separation energies from ${N=92}$ to ${N=93}$.
 Unlike in the cerium isotopic chain, the kink is not sharp but extends to ${N=94}$ forming a prominent bump.
 The gain in energy is about 0.4 MeV, making it one of the strongest changes in two-neutron separation energies over the whole chart of nuclides, away from nuclear shell closures.
-The results call for further studies to elucidate the structure of neutron-rich lanthanum isotopes.
+The results, correlated with a predicted onset of quadrupole deformation for $N\geq 92$, call for further studies to elucidate the structure of neutron-rich lanthanum isotopes.
 Penning-trap mass spectrometry can probe one of the fundamental properties of a nucleus, its mass, with high precision [1, 2].
@@
 The masses of 149,151La are known with a limited precision of 200 and 440 keV [36, 37, 27], respectively, and the mass of 150La is known from a high-precision mass measurement performed at CPT [34].
-In the case of 151Ce93 isotope, the deviation from a smooth trend could be explained by a presence of a long-lived isomer proposed in the ENSDF evaluation [38] but not observed in Ref.
+In the case of 151Ce93 isotope, the deviation from a smooth trend could be explained by the presence of a long-lived isomer proposed in the ENSDF evaluation [38] but not observed in Ref.
 [32].
@@
 We also confirm the sharp kink in the $S_{2n}$ values at $N=93$ for the cerium isotopes.
-The measurement was performed at the Ion Guide Isotope Separator On-Line (IGISOL) facility [65, 66].
-The neutron-rich rare-earth isotopes were produced using $25$-MeV protons impinging on a $15$-mg/cm2-thick ${}^{\text{nat}}$ target.
-The fission fragments, mostly singly charged, were stopped in a helium-filled gas cell (at about 300 mbar) before extracted out through a sextupole ion guide [67].
+The measurements were performed at the Ion Guide Isotope Separator On-Line (IGISOL) facility [65].
+The neutron-rich rare-earth isotopes were produced using a $25$-MeV proton beam with up to 20 $\mu$A of intensity impinging on a $15$-mg/cm2-thick ${}^{\text{nat}}$ target.
+The fission fragments, mostly singly charged, were stopped in a helium-filled gas cell [66, 67] (at about 300 mbar) before being extracted out through a sextupole ion guide [68].
 The ions were further accelerated to $30q$ kV and separated based on their mass-to-charge ratio $m/q$ using a $55$ degree dipole magnet.
@@
 Figure 2 shows two-neutron separation energies for the isotopic chains from caesium (${Z=55}$) to promethium (${Z=61}$) based on AME20 [27], including our new results on 148-153La and 151Ce, as well as the recent CPT measurements for cerium, praseodymium and promethium isotopic chains [35].
-The values were obtained from the mass-excesses as $S_{2n}(Z,N)=\text{ME}(Z,N-2)-\text{ME}(Z,N)+2\text{ME}_{n}$, where $\text{ME}_{n}$ is the neutron mass excess.
+The values were obtained from the mass-excesses as $S_{2n}(Z,N)=\text{ME}(Z,N-2)-\text{ME}(Z,N)+2\text{ME}_{n}$, where $\text{ME}_{n}=8071.3181(4)$ keV is the neutron mass excess [27].
 While the heavier isotopic chains in the rare-earth region exhibit prominent kinks at $N\approx 90$ due to onset of strong quadrupole deformation, the effect seems to vanish when moving to lower proton numbers according to AME20 values.
@@
 Our value for 151Ce confirms the small peak at ${N=93}$ in the cerium chain, coinciding with the transitional region from octupolar to quadrupolar deformation previously studied via $\gamma$-ray spectroscopy.
-For the lanthanum isotopes, our new values reveal a bump about 0.4-MeV in the $S_{2n}$ values at neutron numbers ${N=92-94}$, a unique feature within the neighboring isotopic chains which none of the studied mass models could reproduce.
-The change in the $S_{2n}$ trend agrees with the predicted onset of quadrupole deformation at ${N\geq 92}$ that could be potentially caused by the ${9/2^{+}[404]}$ proton extruder orbital.
-Such a proton extruder could be a plausible explanation for the localisation of the bump within the lanthanum and cerium isotopic chains, as for the other well-known kinks in two-neutron separation energies that span over many isotopic chains, neutron extruder orbitals play the main role.
-More extensive studies on exotic neutron-rich lanthanum isotopes at ${N\geq 90}$ would be highly welcome to elucidate the evolution of nuclear structure in the region.
+Our measurements extend the limits of the known lanthanum masses and reveal around 0.4-MeV bump in the $S_{2n}$ values at neutron numbers ${N=92-94}$, a unique feature within the neighboring isotopic chains not reproduced by any studied mass models.
+The observed change in the $S_{2n}$ trend agrees with the predicted onset of quadrupole deformation at ${N\geq 92}$ that could be potentially caused by the ${9/2^{+}[404]}$ proton extruder orbital.
+This could be a plausible explanation for the localisation of the bump, as analogously neutron extruder orbitals play the main role for the other well-known kinks in two-neutron separation energies.
+Finally, we note that the mass evolution of neutron-rich isotopes further away from stability is crucial for astrophysical calculations, as being one of the key inputs in the determination of neutron-capture reaction rates.
+Specifically, the isotopes reported in this work and the possible structural changes in their vicinity were predicted to be of pivotal importance in the $r$-process [95, 96, 97, 98].
+A better understanding of the nuclear structure in this region is thus of broader interest for the physics community and calls for more extensive experimental and theoretical studies.
 Acknowledgements.